### v2

Abstract
Research surveys have always posed a challenge for beginner researchers who lack of research training. These researchers struggle to understand the directions within their research topic, and the discovery of new research findings within a short time. One way to provide intuitive assistance to beginner researchers is by offering relevant knowledge graphs(KG) and recommending related academic papers. However, existing navigation knowledge graphs primarily rely on keywords in the research field and often fail to present the logical hierarchy among multiple related papers clearly. Moreover, most recommendation systems for academic papers simply rely on high text similarity, which can leave researchers confused as to why a particular article is being recommended. They may lack of grasp important information about the insight connection between ”Issue resolved” and ”Issue finding” that they hope to obtain. To address these issues, this study aims to support research insight surveys for beginner researchers by establishing a hierarchical tree-structured knowledge graph that reflects the inheritance insight of research topics and the relevance insight among the academic papers.
Keywords:
Insight survey Tree-structured Knowledge graph Academic paper Inheritance Relevance
1 Introduction
In recent years, there has been significant progress in the digital environment for academic papers. Databases of academic papers, such as S2ORC[1] and Unarxiv[2], have been continuously improving. The advent of large language models(LLM) like Chatgpt has opened up new possibilities for utilizing these extensive data sources in NLP tasks, including question-answering systems and automatic summarization for academic papers. This advancement enables researchers to conduct research surveys more efficiently using research tools and resources, such as ChatPDF and SciSpace, which have been developed based on ChatGPT and specifically tailored for academic papers understanding[4]. Additionally, knowledge graphs(KG) built on academic papers, such as Connected Papers[3], have also been developed. However, for researchers who are new to a specific academic field or are exploring a research topic, it is difficult to quickly observe, understand, and explore research directions based on large repositories of academic articles[5] solely by simply reading a research paper without a clear target. This inefficient approach to conducting surveys hampers the discovery of research direction and innovative. Therefore, relying solely on paper reading comprehension support software and existing paper recommendation systems may make it difficult to achieve the goal of gaining research insights. Providing an overview of possible research directions and branches within multiple papers on a specific topic may greatly facilitate more efficient exploration of research topics. This process involves the challenging task of conducting in-depth exploration of keywords, theoretical research inheritance, relevance, and their applicability within specific research directions. To address these issue, the objective of this study is to create a knowledge graph for insights survey assistants from multiple academic papers in a specific research topic. It presents a tree structure that shows (1) From the origin of the research task, expand the citation inheritance associations of the research task. (2) Explore the relevance chain among similar research tasks to show relevant research points. It provides potential valuable assistance to researchers in understanding, exploring, and gaining insights into research topics. Additionally, it guides researchers in quickly identifying topic clues and directions for further research expansion. This study contributes to a thorough exploration and visualization of academic topics in the following ways:
(1)Based on the existing academic paper repository, we performed secondary development to fit the requirements of the research insight survey. We integrated a dataset that includes higher-dimensional features, which reflect the citation structure among papers and enable insights content segmentation. (2) To explore the connections of relevance between pairs of papers, we utilize machine learning technology to build a training model for classifying the insight content into ”Issue finding” and ”Issue resolved”. (3) Based on the (1) and (2), we establish a knowledge graph in a hierarchical tree structure. This knowledge graph serves as a guide for beginner researchers exploring research topics to achieve efficient insight survey.
2 Related work
Research survey is a comprehensive investigation conducted to gain an overview of a research field or topic. Especially, it is important for novice researchers to gain a comprehensive understanding of the knowledge structure in their field through research surveys. The development of a KG can be helpful for novice researchers in efficiently understanding and exploring research topics. As previous studies on the KG generated from scientific papers, The PubMed knowledge graph includes meta information such as bio-entities and authors. However, for novice researchers, relying solely on meta information is not sufficient to provide them with enough information about the research content in that field[6]. The knowledge graphs that involve internal information of academic papers include Chan et.al. proposed representation ontology for a four-space integrated KG (background, objectives, solutions, and findings) using NLP technology, as well as Tu et.al.proposed Semantic KG (SKG) that integrates semantic concepts to represent the corpus[7][8].They both classify the content of papers into entity concepts, and associated research key-points in their knowledge graphs. However, the association paths of complex concept maps can be cumbersome, making it difficult for novice researchers to accurately locate academic papers in a specific research direction. This may result in limited insight into the research branch. Jelodar et.al. proposed a semantic framework using probabilistic topic modeling and Gibbs sampling algorithm for a scholarly article recommendation system[9]. However, due to the lack of similarity analysis involving relevance, this makes it difficult for researchers to discover existing issues in a certain research branch. This study aims to expand knowledge mining in the field of insight survey support. It aims to enable researchers to quickly gain insights into research directions. To achieve this, we propose a hierarchical tree structure KG based on the content of academic papers. This graph reflects the inheritance relationships and relevance chain among academic papers.
3 Methodology
3.1 Definition
Research topic: This study focuses on a research topic derived from keywords related to a specific research field. In a research topic, there are multiple research tasks that make up multiple research directions.
Inheritance insight: A survey method that utilizes direct citation relationships between academic papers can be used to identify various branches for exploring inherited relationships in research directions. This study establishes an Inheritance tree to support this kind of insight.
Relevance insight: By analyzing the ”Issue finding” and ”Issue resolved” for each academic paper, a relevance chain of Finding → Resolved → Finding can be established between multiple academic papers to guide researchers in discovering the value of potential direction. This study introduces a relevance tree to facilitate this type of analysis.
The Ideal KG reflecting Inheritance insight and relevance insight survey of ’HotpotQA’ topic are shown in Fig.1-2. The configuration of the KG are shown in Table 1. In Fig.1, the branches p1(Dataset)->p2(Reading Comprehension)->
p4(Evaluation) and p1(Dataset)->p3(Benchmark)->p7(Retrieval) represent two research directions extended by the core paper of the ’HotpotQA dataset’. This graph can help researchers infer the elements of research inheritance in the ’HotpotQA’ topic. For instance, P1 introduced the ’HotpotQA dataset’, P2 utilized this dataset for reading comprehension, and P4 developed evaluation metrics based on reading comprehension. In Figure 2, the path of <Multi-hop branch> p1->Network->p2->Single-hop->p4 and the path of <Retrieval> branch p1->Retrieval->p3->Iteratively->p7 are extensions of the Multi-hop and Retrieval research tasks in the ’HotpotQA’ topic. This tree structure can help researchers infer the correlation factors between subtasks in the ’HotpotQA’ topic. For example, in the ’Multi-hop’ subtask, survey paper P1 provides a list of specific tasks, P2 extends it to the network establishment level, and P3 incorporates single-hop methods.
Table 1: Knowledge graph configuration
Property Content Function
Node_label Paper title
Visually Presenting the theme
of a Research Paper
Node_Title
1. Keyword of paper
2. Issue Resolved in this paper
3. Issue Finding of this paper
Visually display the key
information of research paper,
as a summary.
Edge_direction
1. Inheritance tree
Cite Source –>Cite Destination
2. Relevance tree
Issue Finding –>Issue Resolved
1. Inheritance tree
Reflects the Citation direction
2. Relevance tree
Direction of Relevance chain
Edge_label Co-occurring vocabulary
Helps users understand the
relevance between papers.
Edge_Title
The value of the relevance chain
between papers.
Make it convenient for users
to explore potential directions.
Figure 1: Inheritance tree plan(’HotpotQA topic’)
Figure 2: Relevance tree plan(’HotpotQA topic’)
3.2 Implementation procedure
The overview is shown in Fig.3. The development process is divided into 4 sub-stages: 1. We utilize S2orc as our infrastructure data of KG. S2orc is an academic database that encompasses metadata, bibliographic references, and full text for 8.1 million open access papers[1]. 1. In the first stage, we perform secondary processing based on the S2ORC database to build our insight survey dataset. 2. In the second stage, we use the Sentence Bert model and manually set labels to perform three-class classification (Resolved/Neutral/Finding) on the insight content of each paper to extract the corresponding sentence.3. In the third stage, we use the extracted sentences from phase 2 to perform in-depth analysis on inheritance and relevance chain. We select appropriate papers from the whole insight survey dataset based on certain criteria to generate hierarchical trees.4. In the final stage, we use the pyvis[15] to visualize the tree-structured graph and provide some case studies to demonstrate our findings.
Figure 3: Overview
3.3 Phase1 : Data processing
To build the insight survey dataset, the first step is to select a topic and filter out the relevant sub-dataset from the S2orc database. In this study, we will use ’HotpotQA’ as an example.’HotpotQA’, which is a well-known dataset in the field of NLP. We iterate through each paper in the S2ORC database and extract those that contain the keyword ’HotpotQA’ into our sub-paper library. Our sub-set also includes data-citation information, full-text content, and meta-information of the papers associated with ’HotpotQA’. Next, we extract the text from sections titled ’conclusion’ and ’limitation’ to identify insightful content. We focus on paragraphs with section titles containing ’conclusion’ and ’limitation’ as they address the problems discussed in the research paper and highlight any remaining challenges or limitations. Additionally, we extract relevant citation relationships from the S2orc dataset to create a global network based on the annotations of ’bib-entry’ in the S2orc dataset. The details of insight survey content as described in Table 2.
Table 2: Insight dataset processing (HotpotQA)
Annotation
in S2orc
Description Way of Processing
title Paper title
Use the offset annotation of the paper title
to extract the title text.
section-header
&Paragraph
Section header
To extract paragraphs related to the viewpoints
of ’Finding’ and ’Resolved’ issues, locate a section
header that includes the words ’Conclusion’,
’Discuss’, ’and ’Limitation’. This ensure that the
starting point of the extracted paragraph is appropriate.
bibentry
Link cited
corpusid
Find the corpusid of a paper in S2ORC that matches
the cited paper’s corpusid in the ’HotpotQA’ subset.
3.4 Phase2 : Insight Sentence Extraction
In this section, we discuss the process of insight survey dataset. First, we divide the text in the ’insight-content’ into sentences. We annotate each sentence with the label ”Resolved,” ”Neutral,” or ”Finding” corresponding to the viewpoints the sentence expresses. Next, we use the sentence BERT[10] to vectorize each sentence and adopt SVM classifier to distinguish the corresponding label for each sentence. Finally, sentences with the same ”Resolved” and ”Finding” labels in each article will form the insight sentence of that article.
Sentence Segmentation: Spacy is an open-source natural language processing library for Python that offers an API to access its machine learning trained methods and properties[11]. This work uses the spacy pre-trained model to implement sentence segmentation. Spacy provides a pre-trained English library called en_core_sci_lg which includes a default sentence segmenter [^1] https://allenai.github.io/scispacy/ [^2] https://www.tutorialspoint.com/perform-sentence-segmentation-using-python-spacy. Any complex segmentation patterns that failed were manually fixed.
Human Annotation: For the segmented sentences mentioned above, experts determine the viewpoint of each sentence based on its meaning. The viewpoints include ’Issue Resolved,’ ’Neutral,’ or ’Issue Finding.’ Sentences that thoroughly analyze research methods without explicitly highlighting contributions will be considered ‘Neutral.’ Sentences discussing potential future works in a particular field without specificity will also fall into the ‘Neutral’ category. ’Issue Resolved‘ sentences need at least a combination of contribution, research details, and experimental outcomes, regardless of their positive or negative outcomes. Sentences that ambiguously hint at trends and recommendations may be classified as ‘Issue Finding.’ When categorizing these three viewpoints, the references, tables, and figures of the sentences are unchanged. Table 2 shows specific definitions, distinguishing criteria, and examples of these three labels. The dataset, which consists of insight sentences and labels, has been published on  [^3] https://www.kaggle.com/datasets/dannyleeakira/dataset-for-academic-novelty-insight-survey
Table 3: Sample - Human annotation strategy
Sentence Description Label
In this paper, we present textbrewer, a flexible pytorch-
based distillation toolkit for nlp research and applications.
The target of
this paper
Resolved
Finally, we show that our model facilitates interpretability
by learning an explicit hierarchy…
Things achieved
in this paper
Resolved
For the fourth setting, we pre-train the model for 2 epochs
on the fever dataset, followed by 4 epochs on liar-plus.
Detail of the
method they did
Neutral
We have conducted a series of experiments.
Not related to the
Insight viewpoints
Neutral
Future directions of our work may include using git in
downstream nlp applications where the graph…
Future work of
this paper
Finding
More causal approaches such as amnesiac probing, which
directly intervene in the underlying model’s representations,
may better distinguish between these cases.
Limitation of
this paper
Finding
Training, Sentence Extraction and evaluation: We adpot ’Distilbert-base-uncased’ pre-training model, which was trained using Wikipedia and BookCorpus, for sentence vectorization. We selected Support Vector Machines (SVM) for classifying the vectorized sentences due to their strong generalization performance. We also used 947 labeled sentences for training and validation data. The details of the training data and test data are shown in Table 5. To obtain the optimal SVM parameters, we use the grid search[12] method and find the best parameters to apply to the test. The classification accuracy evaluation is presented in Table 5. There is an imbalance in the number of instances across classes. ”Resolved” class has significantly more instances than the other classes in both the training and test sets. Classification result also shows that ”Resolved” class has higher F1-score as it might be influenced by the larger amount of data. While the ”Neutral” and ”Finding” class has lower F1-scores, indicating challenges in achieving both high precision and recall. Based on the results, we extracted insight sentences of the ”Resolved” and ”Finding” by combining the sentences with corresponding labels within each article.
Table 4: Detail of issue-status dataset
Table 5: Classification result of issue status
Train Test
Resolved 438 142
Neutral 146 53
Finding 126 42
Total 710 237
Precision Recall F1-score
Resolved 0.97 0.81 0.88
Neutral 0.40 0.61 0.48
Finding 0.53 0.68 0.60
Table 5: Classification result of issue status
3.5 Phase3-4 : Hierarchical Tree Construction & Visualization
In this section, we utilize the sentences extracted in phase 2 and the citation information obtained in phase 1 to comprehensively extract the characteristics of the paper. We then employ two strategies to select specific papers in order to construct a tree-structured network.
Similarity calculation: To create the relevance tree, we use the classification results from Phase 2 to determine the elements of the relevance chain (Issue finding → Issue Resolved). We then use embedded insight sentences in sentence transformers[14] to calculate the cosine similarity between insight sentences labeled as ’Finding’ and those labeled as ’Resolved’. Next, we iterate through all the insight sentences in the papers and calculate the relevance chain to generate the relevance matrix. Based on the values in the relevance matrix, we select papers to construct the relevance tree in the following section.
Construction: We extract specific nodes from the insight survey dataset and establish a hierarchical tree-structured according to the following rules, Where N represents the maximum number of root papers(The number of trees in a KG), n is the root sequence of the selected paper, M represents the maximum number of leaves, m is the leaf sequence of the extracted paper, T represents the maximum depth of the tree, and t is the current depth of the tree.
(1) Inheritance tree:
Step 1 - Root node determine:
Sort all the papers in descending order based on the quantity they have been cited. Select the top N papers with the highest citation counts as the root. This operation ensures that the root node has a high level of inheritability throughout the paper library.
Step 2 - Leaf selection: From the candidate group of papers that cite the root node paper, select the top m papers with the highest citation counts as the leaf nodes of rootn.
Step 3 - Parent node update: Set each leaf node as a new root node and repeat Step 2. If a root node is not cited (unable to generate a leaf), the branch is terminated.
(2) Relevance tree:
Step 1 - Root node determine:
Calculate the average of similarity score for each paper in the relevance matrix. Select the top N papers with the highest similarity score as the root nodes. This operation ensures that the root node has a high correlation index throughout the paper library.
Step 2 - Leaf selection: From the candidate group of papers that have relevance chain with the root node paper. Select the top m papers with the highest similarity score as the leaf nodes of root n.
Step 3 - Parent node update: Set each leaf node as a new parent node and repeat Step 2. If a root node is not have relevance chain (unable to generate a leaf), the branch is terminated.
(3) Common Rules:
Rule 1:
The selected nodes cannot be selected again.
Rule 2: Repeat Steps 1-3 until reaching the upper limit of N or the maximum depth limit value T. The network is then generated.
To provide researchers with a visual understanding(Fig.1-Fig.2) of the guidance on research directions provided by these selected papers. We use pyvis to visualize KG based on the configuration of Table 1. In addition to the co-occurring vocabulary of insight sentences between papers, we also extract keywords specific to each paper to capture its unique concept. When selecting keywords for each paper in the KG, we combine the ‘insight content’ from all papers in the subtree to create the overall text. The ’insight content’ of a specific article within the subtree is used as the partial text. Tf-idf[13] method used to extract partial text keywords from the global text composed of multiple paper-texts in a tree, and then display them as paper keywords in the node titles, and a part of the representative sample branch included insight sentences is shown in Fig.LABEL:figure3.
4 Conclusion
This study developed two type of hierarchical tree-structured knowledge graph called inheritance tree and relevance tree to support insight survey for research beginners. The process consists of four stages: data processing, insight sentence extraction, hierarchical tree construction, and KG visualization. First, we conducted high-dimensional secondary development of the S2ORC dataset to create an insight survey dataset that includes citation information and insight content. Then, using machine learning techniques, we extracted sentences from the insight survey dataset that express insight viewpoints on ’Issue finding’ and ’Issue resolved’. These sentences were parsed and used to construct a relevance matrix. Finally, based on the citation information and relevance matrix, we generated two types of hierarchical tree structures: Inheritance tree and Relevance tree. The generated KG demonstrate their rationality, indicating that they can provide key-information to assist researchers in gaining insights into the directions of the research topic. The knowledge graph also exhibits interpretability and potential for further development. For the future expansion and improvement of this study, the following points are proposed:
(1) Incorporate Additional Text for Text Similarity and relevance Chain Computation: Currently, this study only utilizes the text from the ‘conclusion’ and ‘limitation’ section. However, the Abstract and Introduction sections, specifically the part discussing previous issues, contain novel elements related to previous research topics and the subject of this study. Therefore, integrating these texts for high-dimensional training can optimize the content in knowledge graph.
(2) We conducted an objective evaluation of the classification accuracy of the ’Issue finding’ and ’Issue resolved’ viewpoints. However, integrating subjective evaluations from researchers into the generated knowledge graph remains a future challenge.